# url]https://person.dibris.unige.it/ricca-filippo/

### v1

@@ 4.3 RQ3 (Tools and Platforms)
Table 4 shows that we identified in total 100 tools that support the TA solutions strategies (RQ2) and that were discussed in the analyzed documents in the three iterations. Among these tools, the table lists the twenty-two most frequently cited and discussed ones. Considering the third iteration, the top-5 tools are: Applitools, [^1] https://applitools.com Testim, [^2] https://www.testim.io Functionize, [^3] https://www.functionize.com AccelQ, [^4] https://www.accelq.com and Mabl. [^5] https://www.mabl.com
@@
4.4 RQ4 (Generalizability)
Table 2, Table 3, and Table 4 report TA problems, solutions, and existing tools/platforms that we have identified in the three iterations of the conducted grey literature analysis: first [47], second [48] (in green), and third iteration (in light blue) documented for the first time in this work. In the first iteration, we analyzed 156 documents (see Table 1), in the second one we additionally considered 95 documents (+60.8% to the initial set of 156 documents of [47]), and in this third iteration, we additionally considered 91 documents (+36.2% concerning the total number of documents, that is, 251 found in the first two iterations).
The three different iterations allow us to: (i) built a taxonomy of TA problems composed of 47 subcategories (40 after the second iteration, and 35 after the first one), grouped into 7 main categories (6 after the first and the second iterations), and in which 917 individual occurrences have been distributed (545 after the second iteration, and 423 after the first one); (ii) built a taxonomy of TA solutions composed of 59 subcategories (44 after the second iteration, and 35 after the first one), grouped into 8 main categories (5 after the second iteration, and 4 after the first one), and in which 972 individual occurrences have been distributed (607 after the second iteration, and 466 after the first one); and (iii) identified a list of 100 tools and platforms (67 after the second iteration, and 50 after the first one).
With respect to the two initial taxonomies, in detail, we have only partially extended the taxonomy related to TA problems, leaving it almost stable with respect to the one defined initially in [47]. In fact, during the second iteration, we only introduced 5 new subcategories of TA problems, resulting in a total of 44 individual occurrences out of 545. In the third iteration, we added 7 additional subcategories, which accounted for 25 individual occurrences out of 917. We also made only partial progress in extending the taxonomy related to AI-based solutions. Specifically, during the second iteration, we introduced one new category (i.e., Test Execution) and 9 subcategories, which accounted for 80 individual occurrences out of 607. In the third iteration, we added 3 additional categories (i.e., Test Optimization, Test Process, and Test Type) along with 15 subcategories, resulting in 77 individual occurrences out of 972.
Regarding tools and platforms, however, the situation is completely different. In a short time, their number has grown significantly, and new options have appeared on the market. Among the new tools and platforms, we find ChatGPT, Aqua, TestRigor, and pCloudy. ChatGPT, an advanced language model from OpenAI that generates human-like text, is primarily used for creating test scripts in software testing and was the most frequently mentioned tool in the documents we reviewed, with 14 total occurrences. Following that, we have Aqua, TestRigor, and pCloudy. Aqua is a test management platform that leverages AI to optimize test planning, execution, and tracking. Offering cloud-based solutions for both manual and automated testing, it integrates with various tools to improve testing efficiency and effectiveness. TestRigor is an automated testing tool designed for creating and running end-to-end tests using plain English commands. It simplifies test creation and maintenance. pCloudy is a unified app testing platform that ensures app quality across various devices and browsers. It offers cloud-based manual and automated testing with access to thousands of real devices. Leveraging AI, pCloudy enhances integration, efficiency, and effectiveness in testing
In conclusion, regarding the generalizability of the taxonomies, we can state that in our analysis, we have only partially extended the initial taxonomies. The list of problems and solutions has remained largely consistent with the previous versions, with only a few minor exceptions. On the other hand, the list of tools has shown significant and consistent growth across the iterations, doubling their number from the first to the third iteration.
@@ 4.4 RQ4 (Generalizability)
4.5 RQ5 (Problems vs. Solutions)
Figure 3 utilizes a Sankey diagram to illustrate the relationships between problems and AI-based solutions in the test planning phase: problems are shown on the left, while solutions are shown on the right. In the planning phase, the most relevant problem (the larger node on the left) identified by the community is the identification of test paths, i.e., a specific sequence of actions that are executed to verify that a particular feature or functionality of the application under test is working as expected. This sequence can involve clicking on links, filling out forms, submitting data, and interacting with various elements on the web pages. In particular, testers are interested in identifying the critical paths in the app to test them thoroughly. In terms of solutions adopted, automatic test generation based on crawling and user behaviors are the most frequently discussed. However, as shown in the diagram, other solutions for identifying critical paths are: predict faulty areas in the application under test and intelligent test analytics. It is worth noting that test process management has become an increasingly relevant issue over the years. In fact, there has been a noticeable rise in individual discussions on this topic within the community. As shown in Table 2, there were no occurrences of this issue identified in the first iteration, eight occurrences in the second iteration, and twenty occurrences in the most recent iteration. The most commonly used solution for this problem is using AI and data analytics to guide the software testing lifecycle. This approach leverages various technologies and methods to make informed decisions based on data rather than relying solely on manual processes or intuition.
@@
Figure 5 represents the Sankey diagram problems/solutions in the authoring phase and shows that the most relevant problems in this phase are related to the manual effort needed to develop/implement test cases in test scripts and creating effective test data (see the two larger nodes on the left). Among the solutions to limit the manual effort in implementing test cases, we find that the predominant category, as we would expect, is automatic generation. The subcategories include all the automatic generation methods found, such as using screen mockups as a starting point, employing a web crawler, and even utilizing generative AI tools/platforms. Regarding the problem of generating effective test data, the main solution is to rely on AI-based tools capable of generating it. It is interesting to note that test object identification—the ability of testing tools/platforms to correctly recognize and interact with various elements (objects) on a web page, e.g. dynamic or modal elements, and cross-platform problems—cross-platform testing involves verifying that a software application functions correctly across different environments, including various operating systems, browsers, and devices. Planning cross-platform testing presents unique challenges due to the diversity and complexity of environments on which the application must be tested—remain among the most relevant problems, in all the three considered iterations. In particular, the identification of test objects is faced with several strategies such as (i) object recognition engines (i.e., identification of testing elements in the GUI); (ii) robust element localization (e.g., use of multiple sources to localize elements in the GUI); (iii) visual testing (i.e., automated visual checks of the GUI by using of computer vision); and (iv) intelligent fault prediction strategies.
@@
5 Discussion
Observations and Evidences. By analyzing the results collected in the three iterations of our study, we derived the following observations and evidence.
  • O1:
    We can identify TA problems for which AI-enhanced existing solutions seem to be promising and largely studied. In particular, we observed that test authoring (i.e., test creation) is the most investigated TA problems. We identified 34 tools (e.g, Applitools, Functionize, Testim, Testrigor) that use AI-based technology for automatically generating test cases using different technologies, such as NLP techniques or GUI-based capture and reply techniques to automatically generate test scripts, aiming at limiting the manual human intervention. O1 is of interest, in particular, for practitioners who could easily identify the problems better supported by existing tools, and also for researchers to identify unexplored problems for which propose innovative solutions.
  • O2:
    We can identify TA problems that could be faced by AI-enhanced existing solutions implemented in AI-based tools. For instance, a well-known challenge in GUI testing is the creation of effective oracles [35]. The use of oracles based on visual testing, using AI and computer vision approaches, is suggested in the grey literature as one of the possible ways to face this challenge. Furthermore, 13 tools that support oracle visual testing have been identified, e.g., Applitools, AI Testbot, Mabl, Sealights, Testim, and Test.ai. O2 is of interest, in particular, for practitioners for quickly selecting the most appropriate solutions for their TA problems and the most adequate tools that support the solutions to their TA problems.
  • O3:
    We can identify TA problems that require solutions involving several phases of TA. For instance, in the test planning phase, we defined problems such as planning what to test, identifying critical paths in the application under test, and managing the entire test process, among others. We observe that the latter problem raised attention only recently (in the third iteration of our study), and it is faced by mainly applying AI data-driven decision strategies. The second problem is mainly faced with automatic test creation solutions, ranging from crawling the application under test to the creation of test scripts using user behaviors. The first problem instead is faced by solutions involving test script creation (e.g., automatic test creations by focusing on application areas that are predicted as more buggy), test selection (e.g., based on fault prediction), test execution (e.g., adoption of intelligent test re-execution strategies), and, finally, test debugging approaches (e.g., adoption of test analytics). O3 is of interest, in particular, for researchers that can better highlight as some TA problems are addressed from different perspectives, i.e., for some problems, specific ad-hoc solutions can be adequate while, for other problems, more complex solutions need to be studied.
  • O4:
    We identified solutions presented in the grey literature that are not supported by existing available tools. For instance, nevertheless, test generation with mockups and dynamic properties identified from observed user behaviors are listed among possible solutions supported by AI, for automatically generating tests, we did not identify any existing tools that provided such capabilities. However, it is important to keep in mind that the absence of a tool does not necessarily indicate a lack of existing solutions for a particular problem. It could simply mean that the specific tooling solutions were not mentioned in the literature due to the incompleteness of our analysis. For instance, decoupling the test framework from the host environment is referenced, in the grey literature, as one possible solution for facilitating cross-platform testing. However, it seems that this solution is not adequately supported by the existing AI-enhanced tools, that mainly provide approaches that allow the identification of different environmental configurations, to face cross-platform testing. Another solution that seems to be not adequately supported by tools concerns the execution of test cases with mock responses: no tools support the construction of mock objects that can be used in TA. Concerning the test selection and optimization, solutions aiming at prioritizing test cases, removing unnecessary test cases and GUI-based testing seem to be not adequately supported by existing AI-enhanced tools. O4 is of interest to both professionals and researchers to develop innovative tools and technologies capable of supporting the identified solutions.
  • O5:
    In general, we identified tools that support solutions across various phases of test automation and address multiple test automation problems. However, there are also, though less common, tools that are specific to a particular phase or problem. For instance, tools such as Applitools, Functionize, and Mabl are able to support different TA phases, e.g., test creation, maintenance, execution. Conversely, tools such as Appium, Testsigma, and TestComplete seem to be more specific, thus mainly supporting a given testing phase, e.g., Appium and Testsigma focus on test creation, while TestComplete focuses on visual testing. O5 is of interest for practitioners for selecting the most appropriate tools to use in their business, by taking into account the problems they have to face and also other aspects such as specificity and flexibility of tools.
5.1 Threats to validity
@@ 6 Developers Interviews
To get a more complete picture, we have conducted semi-structured interviews with five researchers/practitioners with various backgrounds and levels of expertise, focusing on the types of problems and solutions found in our taxonomies.
@@ 6.1 Interviews Design
We selected five developers/researchers from personal contacts using convenience sampling and conducted semi-structured interviews to understand whether they perceive AI as a valuable asset that aligns with their needs and objectives in their testing activities.
Table 5 provides details about the interviewees, including their position, expertise, and years of experience. Among the interviewed candidates the lowest value is 6 years and the highest is 20 years (median=14). Our pool includes professionals involved in strategic decision-making, as well as technical personnel and a researcher.
@@
All interviews were conducted remotely via Teams video calls, each lasting approximately 60 minutes. The interviews were transcribed using Teams’ automated speech recognition tool, which converts audio/video files into text. After generating the automated transcriptions, we reviewed and manually corrected transcription errors. Finally, we proceeded with open coding of the transcribed interviews in which different parts of the transcribed text were tagged.
@@ 6.2.1 Problems
In end-to-end testing, the interviewees mentioned that a primary challenge is testability. The lengthy and complex nature of E2E test cases makes testing at the API level very difficult, especially without a clear specification. E2E tests are often fragile due to the lack of an API, the absence of static checking, synchronization issues, and missing types. Creating a model for these tests is considered a valuable investment, as it can result in more maintainable and robust tests, especially if the web app model creation and maintenance can also be automated.
@@ 6.2.2 Solutions
The participants emphasized the use of AI for identifying dependencies between services and detecting obsolete test cases due to the lack of domain knowledge in complex codebases. Another proposed solution is to implement anomaly detection using test execution logs to monitor the system and production data to help anticipate failures before they occur. An interesting case was the usage of LLMs for mutation testing: when new test cases were introduced in the test suite, the web app was modified with artificially generated faults to assess the quality of the evolved test suite.
@@
Ultimately, the participants see AI as a collaborative partner in testing, code review, and pair programming. With accountability resting on the developers, they write the tests (serving as specifications), while the AI generates the code. The final message was to use AI to generate boilerplate code while keeping decision-making a human task.
@@ 6.2.3 Tools
The interviewees described employing a diverse collection of AI algorithms in their work. These range from traditional machine learning algorithms, such as Support Vector Machines (SVM) and decision trees. An interview mentioned “conventional AI” strategies such as search-based optimization algorithms, model-based approaches such as user interaction flows, or Markov Chain Monte Carlo for prioritizing tests and tracking the frequency of test changes. The participants also mentioned the usage of computer vision techniques like object recognition to find web elements that are difficult to find in the DOM, internationalization (I18n) testing, or automating the visual oracle in tests.
@@
Our participants highlighted their reliance on in-house tools and existing frameworks, over custom code bases that are hard to adapt to each specific customer’s need. They also noted the integration of advanced generative AI models, such as OpenAI’s ChatGPT and Microsoft’s Copilot, into their daily activities for code generation and review.
6.2.4 Open Challenges of AIaTA

### v2

@@ 4.3 RQ3 (Tools and Platforms)
Table 6 shows that we identified in total 100 tools that support the TA solutions strategies (RQ2) and that were discussed in the analyzed documents in the three iterations (67 after the second iteration, and 50 after the first one). Among these tools, the table lists the twenty-two most frequently cited and discussed ones. Considering the third iteration, the top-5 tools are: Applitools, [^1] https://applitools.com Testim, [^2] https://www.testim.io Functionize, [^3] https://www.functionize.com AccelQ, [^4] https://www.accelq.com and Mabl. [^5] https://www.mabl.com
@@
Our survey extension highlighted a significant increase in the availability of new tools for developers. Their number has grown significantly in a short time, and new options have appeared on the market. We find ChatGPT, Aqua, TestRigor, and pCloudy among the new tools and platforms. ChatGPT, an advanced language model from OpenAI that generates human-like text, is primarily used for creating test scripts in software testing and was the most frequently mentioned tool in the documents we reviewed, with 14 total occurrences. Following that, we have Aqua, TestRigor, and pCloudy. Aqua is a test management platform that leverages AI to optimize test planning, execution, and tracking. Offering cloud-based solutions for both manual and automated testing, it integrates with various tools to improve testing efficiency and effectiveness. TestRigor is an automated testing tool designed for creating and running end-to-end tests using plain English commands. It simplifies test creation and maintenance. pCloudy is a unified app testing platform that ensures app quality across various devices and browsers. It offers cloud-based manual and automated testing with access to thousands of real devices. Leveraging AI, pCloudy enhances integration, efficiency, and effectiveness in testing.
@@
4.4 RQ4 (Problems vs. Solutions)
Figure 3 utilizes a Sankey diagram to illustrate the relationships between problems and AI-based solutions in the test planning phase: problems are shown on the left, while solutions are shown on the right. In the planning phase, the most relevant problem (the larger node on the left) identified by the community is the identification of test paths, i.e., a specific sequence of actions that are executed to verify that a particular feature or functionality of the application under test is working as expected. This sequence can involve clicking on links, filling out forms, submitting data, and interacting with various elements on the web pages. In particular, testers are interested in identifying the critical paths in the app to test them thoroughly. In terms of solutions adopted, automatic test generation based on crawling and user behaviors are the most frequently discussed. However, as shown in the diagram, other solutions for identifying critical paths are predicting faulty areas in the application under test and intelligent test analytics. It is worth noting that test process management has become an increasingly relevant issue over the years. In fact, there has been a noticeable rise in individual discussions on this topic within the community. As shown in Table 4, there were no occurrences of this issue identified in the first iteration, eight occurrences in the second iteration, and twenty occurrences in the most recent iteration. The most commonly used solution for this problem is using AI and data analytics to guide the software testing lifecycle. This approach leverages various technologies and methods to make informed decisions based on data rather than relying solely on manual processes or intuition.
@@
Figure 5 represents the Sankey diagram problems/solutions in the authoring phase and shows that the most relevant problems in this phase are related to the manual effort needed to develop/implement test cases in test scripts and create effective test data (see the two larger nodes on the left). Among the solutions to limit the manual effort in implementing test cases, we find that the predominant category, as we would expect, is automatic generation. The subcategories include all the automatic generation methods found, such as using screen mockups as a starting point, employing a web crawler, and even utilizing generative AI tools/platforms. Regarding the problem of generating effective test data, the main solution is to rely on AI-based tools capable of generating it. It is interesting to note that test object identification—the ability of testing tools/platforms to correctly recognize and interact with various elements (objects) on a web page, e.g. dynamic or modal elements, and cross-platform problems—cross-platform testing involves verifying that a software application functions correctly across different environments, including various operating systems, browsers, and devices. Planning cross-platform testing presents unique challenges due to the diversity and complexity of environments on which the application must be tested—remain among the most relevant problems, in all the three considered iterations. In particular, the identification of test objects is faced with several strategies such as (i) object recognition engines (i.e., identification of testing elements in the GUI); (ii) robust element localization (e.g., use of multiple sources to localize elements in the GUI); (iii) visual testing (i.e., automated visual checks of the GUI by using of computer vision); and (iv) intelligent fault prediction strategies.
@@
Figure 10: Representation of the self-healing mechanism on the Login example.
5 Qualitative Examples: Automated Generation and Self-healing Test Scripts
In this section, we exemplify three of the most popular AI-based TA features, aiming to provide a more concrete understanding of some of the AI solutions proposed in Table 5: codeless test script generation, automated test script generation using GenAI and self-healing test scripts. We illustrate these features in the web domain using testRigor, [^6] https://testrigor.com/ one of the most popular tools identified in the third iteration (see Table 6). We selected testRigor because it offered a 14-day trial license for full use of the tool (although limited to a single test suite). It is important to emphasize that similar features are also implemented, perhaps in different ways, in other tools identified by our taxonomy.
testRigor falls into the category of codeless testing tools. It enables the creation of test scripts using plain English commands, eliminating the need for traditional programming. By leveraging advanced NLP algorithms, testRigor translates these plain English instructions into executable test scripts and executes them. testRigor is capable of interpreting and translating straightforward commands in test scripts, such as "open" to launch a webpage in a browser, "click" to interact with a button, "insert" to input text into a field, and "check" to validate the presence of specific text on a webpage. For example, a simple testRigor test to verify the successful log-in on a web application could look like this:
enter "student" into "Username"
enter "Password123" into "Password"
click "Submit"
check if page contains "You successfully logged in!"
In addition to being codeless, the testRigor tool leverages GenAI to automatically generate NLP-based test scripts from informal natural language descriptions of a test case, simplifying the test development process. For instance, a description like "Test checkout process" enables the AI to generate a relevant test script description, reducing the need for manual input. Similarly, a more detailed description, such as "Develop a full test for adding a MacPro to a shopping cart", for a given e-commerce web app, allows testRigor to produce the following complete test script:
Click on "search bar" at the top of the page
Enter "MacBook" into "search bar"
Press Enter
Click on the first hyperlink containing "MacBook"
Click on "Add to Cart"
Click on "Go to Cart"
Check that page contains "MacBook"
Check that page contains "Quantity: 1"
Although this approach minimizes manual effort, its accuracy and usefulness depend on the clarity and specificity of the descriptions provided.
In addition to its generation features, testRigor also offers the ability to enable self-healing test scripts. Self-healing in software engineering refers to systems that automatically detect and fix errors without human intervention. Through advanced algorithms and monitoring, these systems identify issues, take corrective actions, and restore functionality, minimizing downtime and improving reliability. In test automation, self-healing scripts can adapt to changes in the application under test (AUT) by automatically correcting themselves when a test fails.
Imagine an automated test script designed to interact with a button on a web page or software interface by identifying it through a specific ID locator. This method functions seamlessly until a new version of the software alters the button’s ID. As a result, the test fails, unable to locate and interact with the button. Such failures lead to maintenance challenges that require developers or testers to update the test scripts with the new locator details. This process is often tedious and resource-intensive. Self-healing test automation addresses this issue by introducing adaptability directly into the testing framework. Instead of failing outright when encountering a modified element, such as a button with a changed ID, a self-healing system actively attempts to identify the element using alternative methods, such as analyzing attributes, context, and surrounding elements or using computer vision algorithms. This capability allows the test to continue running without human intervention, minimizing disruptions and significantly reducing the effort required for test maintenance.
testRigor implements self-healing tests by leveraging a component called Vision AI to adapt to specification changes for rules and individual commands. Instead of failing, testRigor analyzes the screen to find alternative ways to achieve the intended actions. Vision AI uses computer vision to enable machines to interpret and understand visual elements of user interfaces—such as images, icons, buttons, and text—just like a human tester. This approach allows the test script to quickly adjust to changes in a web app, making it particularly effective for testing complex graphical user interfaces. To better understand the concept of self-healing test scripts and how this mechanism is implemented in testRigor, let us consider a web page containing a Login form (version n), as shown above in Figure 10. Since testRigor operates as a human tester would, we created a test script to perform the Sign-In action using the UI text as locators, as already shown above in the first test script example. Now, let us suppose that the application under test has been updated (version n+1), and the button web element has been changed to ’Submit’ (as shown below in Figure 10).
Figure 11: Test script execution with the self-healing mechanism enabled (represented by the box marked in orange).
A standard, non-self-healing test script would fail in this scenario because it would not locate the ’Login’ text on the button. However, when executed with testRigor, which incorporates a self-healing feature, the test script successfully adapts and does not fail. As illustrated in Figure 11, the execution results show that the first step and the third step—corresponding to opening the webpage and executing the check command after clicking the Login button—are marked in green, indicating successful completion. Step two, however, is highlighted in orange, signaling that the self-healing mechanism was activated after entering the username and the password, to identify and adjust for the modified button in the updated version. Additionally, a link labeled “More details…” provides access to a detailed summary of the adjustments made. The test script execution demonstrates that when UI changes occur, testRigor’s AI attempts to identify alternative ways to interact with the affected element. By clicking the “More details…” button, a popup appears displaying the original test script alongside the AI-modified version with updated test steps. If reverting to the previous version is preferred, the “Rollback to this” option can be selected. Alternatively, if the AI-generated modifications are satisfactory, they can be retained as the final version.
@@ 6 Developers Interviews
To get a more complete picture, we have conducted semi-structured interviews with researchers and practitioners with various backgrounds and levels of expertise, focusing on the types of problems and solutions found in our taxonomies.
@@ 6.1 Interviews Design
The study focused on software engineers with experience in software testing and artificial intelligence. Convenience sampling was utilized to recruit participants, allowing for efficient enrollment within the study’s constraints. However, this approach inherently carries a risk of bias due to the non-random selection process. To address this limitation, we encouraged word-of-mouth promotion of the interview opportunity, leading to the recommendation of diverse and appropriate candidates. Only individuals with a minimum of five years of professional experience in software engineering were included. This criterion ensured that participants had sufficient exposure to both software testing and AI practices. Efforts were made to select participants representing a range of ages, genders, and professional roles. This approach aimed to incorporate diverse perspectives and reduce the risk of homogeneity in viewpoints. Ultimately, five developers/researchers agreed to participate in our study and conducted semi-structured interviews to understand whether they perceive AI as a valuable asset that aligns with their needs and objectives in their testing activities.
Table 7 provides details about the interviewees, including their position, expertise, and years of experience. Among the interviewed candidates the lowest value is 6 years and the highest is 20 years (median=14). Our pool includes professionals from a company in Sweden involved in strategic decision-making, as well as technical personnel and a researcher at an applied science university in Austria with substantial experience in the industry.
@@
All interviews were conducted remotely via Teams video calls, each lasting approximately 60 minutes. The interviews were transcribed using Teams’ automated speech recognition tool, which converts audio/video files into text. After generating the automated transcriptions, we reviewed and manually corrected transcription errors. Finally, we proceeded with open coding of the transcribed interviews in which different parts of the transcribed text were tagged. The sanitized text is available in our replication package [53].
@@ 6.2.1 Problems
The usage of AI touched on several aspects of the software development process. For example, it is used for requirement elicitation during interviews with clients and customers to complement information and write summary reports.
@@
In end-to-end testing, the interviewees mentioned that a primary challenge is testability. The lengthy and complex nature of E2E test cases makes testing at the API level difficult, especially without a clear specification. E2E tests are often fragile due to the lack of a clear API, static checking, synchronization issues, and missing types. Creating a model for these tests is considered a valuable investment, as it can result in more maintainable and robust tests, especially if the web app model creation and maintenance can also be automated.
@@ 6.2.2 Solutions
The participants emphasized the use of AI to identify dependencies between services and detect obsolete test cases due to the lack of domain knowledge in complex codebases. Another proposed solution is to implement anomaly detection using test execution logs to monitor the system and production data to help anticipate failures before they occur. An interesting case was the usage of LLMs for mutation testing: when new test cases were introduced in the test suite, the web app was modified with artificially generated faults to assess the quality of the evolved test suite.
@@
Ultimately, the participants see AI as a collaborative partner in testing, code review, and pair programming. With accountability resting on the developers, they write the tests (serving as specifications), while the AI generates the code. The usage of “AI to create test cases” was reported to be “in the very early stages at the moment”. It was mentioned that AI was used to create API test mockups that developers complement manual testing by adding logic.
The final message was to use AI to generate boilerplate code while keeping decision-making a human task.
@@ 6.2.3 Tools
Our participants highlighted their reliance on in-house tools and existing frameworks like Microsoft Azure DevOps, over individual tools or custom code bases that are hard to adapt to each specific customer’s need. They also noted the integration of advanced generative AI models, such as OpenAI’s ChatGPT and Microsoft’s Copilot, into their daily activities for code generation and review.
The interviewees described employing a diverse collection of AI algorithms in their work. These range from traditional machine learning algorithms, such as Support Vector Machines (SVM) and decision trees. An interview mentioned “conventional AI” strategies such as search-based optimization algorithms, model-based approaches such as user interaction flows, or Markov Chain Monte Carlo for prioritizing tests and tracking the frequency of test changes. The participants also mentioned the usage of computer vision techniques like object recognition to find web elements that are difficult to find in the DOM, internationalization (i18n) testing, or automating the visual oracle in tests.
7 Threats to Validity
Another potential threat to validity pertains to the interviews and the open coding process. While these were conducted solely by the third author, his prior experience with similar studies and analyses [22] lends credibility to the approach. The structured interviews were limited to five participants, a constraint largely due to the difficulty of engaging developers actively engaged in their professional responsibilities. Despite the small sample size, it is important to highlight that all participants were senior developers. This ensures a solid foundation for the insights gathered, given the depth of experience and expertise they brought to the discussions. Although larger or more diverse samples could potentially lead to different conclusions, our analysis of the open-ended responses revealed recurring themes and no emergence of new ones, suggesting that saturation was reached. Consequently, we believe that, despite the limitations, the sample size is adequate for this study. The use of convenience sampling could lead to a lack of representativeness in the sample, potentially skewing findings. To mitigate this, efforts were made to diversify the sample as much as possible within the method’s constraints. However, it is acknowledged that the findings may not be fully generalizable to the broader population of software engineers.
8 Discussion
8.1 Observations from the Survey
8.1.1 Generalizability
In the first iteration, we analyzed 156 documents (see Table 1), in the second one we additionally considered 95 documents (+60.8% to the initial set of 156 documents of [55]), and in this third iteration, we additionally considered 91 documents (+36.2% concerning the total number of documents, that is, 251 found in the first two iterations). Concerning the two initial taxonomies, in detail, we have only partially extended the taxonomy related to TA problems, keeping it largely unchanged from the original version [55]. In fact, during the second iteration, we only introduced 5 new subcategories of TA problems, resulting in a total of 44 individual occurrences out of 545. In the third iteration, we added 7 additional subcategories, which accounted for 25 individual occurrences out of 917. We also made only partial progress in extending the taxonomy related to AI-based solutions. Specifically, during the second iteration, we introduced one new category (i.e., Test Execution) and 9 subcategories, which accounted for 80 individual occurrences out of 607. In the third iteration, we added 3 additional categories (i.e., Test Optimization, Test Process, and Test Type) along with 15 subcategories, resulting in 77 individual occurrences out of 972. In conclusion, regarding the generalizability of the taxonomies, we can state that in our analysis, the list of problems and solutions has remained largely consistent with the previous versions, with only a few minor additions. On the other hand, the list of tools has shown significant and consistent growth across the iterations, doubling their number from the first to the third iteration.
8.1.2 Prevalence of Web E2E Testing
Our search query was designed to encompass software testing broadly, without focusing on any specific type of system. Nevertheless, our survey of the grey literature revealed that most articles centered on E2E testing for mobile and web applications, followed by unit and API testing.
8.1.3 AI for Innovative Test Authoring
We identified TA problems for which AI-enhanced existing solutions seem to be promising and largely studied. In particular, we observed that test authoring (i.e., test creation) is the most investigated TA problem. We identified 34 tools (e.g, Applitools, Functionize, Testim, Testrigor) that use AI-based technology for automatically generating test cases using different technologies, such as NLP techniques or GUI-based capture and reply techniques to generate test scripts, aiming at limiting the manual human intervention. Research approaches have been proposed to this aim [4, 12]. Practitioners can use our taxonomy to easily identify the problems better supported by existing tools whereas researchers can identify unexplored problems for which they can propose innovative solutions.
8.1.4 AI-assisted Visual Oracle
We identified TA problems that could be faced by AI-enhanced existing solutions implemented in AI-based tools. For instance, a well-known challenge in GUI testing is the creation of effective oracles [43, 24, 25]. The use of oracles based on visual testing, using AI and computer vision approaches, is suggested in the grey literature as one of the possible ways to face this challenge. This aligns with recent research visual approach to web testing [66, 71]. Furthermore, 13 tools that support oracle visual testing have been identified, e.g., Applitools, AI Testbot, Mabl, Sealights, Testim, and Test.ai. Practitioners can use our findings to quickly select the most appropriate solutions for their TA problems and the most adequate tools that support the solutions to their TA problems.
8.1.5 Phase-specific and Multi-Phase Solutions for Smarter Test Automation
In general, we identified tools that support solutions across various phases of test automation and address multiple test automation problems. However, there are also, though less common, tools that are specific to a particular phase or problem. For instance, tools such as Applitools, Functionize, and Mabl can support different TA phases, e.g., test creation, maintenance, and execution. Conversely, tools such as Appium, Testsigma, and TestComplete seem to be more specific, thus mainly supporting a given testing phase, e.g., Appium and Testsigma focus on test creation, while TestComplete focuses on visual testing. Our findings are of interest to practitioners in selecting the most appropriate tools to use in their business, by taking into account the problems they have to face and also other aspects such as the specificity and flexibility of tools.
We also identified TA problems that require solutions involving several phases of TA. For instance, in the test planning phase, we defined problems such as planning what to test, identifying critical paths in the application under test, and managing the entire test process, among others. We observe that the latter problem raised attention only recently (in the third iteration of our study), and it is faced by mainly applying AI data-driven decision strategies [14]. The second problem is mainly faced with automatic test creation solutions [3], ranging from crawling the application under test [4] to the creation of test scripts using user behaviors [12]. The first problem instead is faced by solutions involving test script creation (e.g., automatic test creations by focusing on application areas that are predicted as more buggy), test selection (e.g., based on fault prediction), test execution (e.g., adoption of intelligent test re-execution strategies), and, finally, test debugging approaches (e.g., adoption of test analytics). Researchers can use our findings to better highlight as some TA problems are addressed from different perspectives, i.e., for some problems, specific ad-hoc solutions can be adequate while, for other problems, more complex solutions need to be studied.
8.1.6 Open Challenges for AIaTA
We identified solutions presented in the grey literature that are not supported by existing available tools. For instance, nevertheless, test generation with mockups and dynamic properties identified from observed user behaviors are listed among possible solutions supported by AI, for automatically generating tests, we did not identify any existing tools that provided such capabilities. Non-AI research tooling has been proposed in literature [10], but AI-driven solutions remain limited, with only a few proposals, such as those leveraging deep learning [59], introduced to date. However, it is important to keep in mind that the absence of a tool does not necessarily indicate a lack of existing solutions for a particular problem. It could simply mean that the specific tooling solutions were not mentioned in the literature due to the incompleteness of our analysis. For instance, decoupling the test framework from the host environment is referenced, in the grey literature, as one possible solution for facilitating cross-platform testing.
It seems that this solution is not adequately supported by the existing AI-enhanced tools, that mainly provide approaches that allow the identification of different environmental configurations, to face cross-platform testing. Conversely, this issue has been addressed in the white literature for web [77] and mobile domains [77] with approaches based on robotic arms [75], deep reinforcement learning [79], and computer vision to enable the Chouette Crawler to help accelerate the mobile and web app testing cycle at Duolingo [73].
Another solution that seems to be not adequately supported by tools concerns the execution of test cases with mock responses: no tools support the construction of mock objects that can be used in TA. Concerning the test selection and optimization, solutions aiming at prioritizing test cases, removing unnecessary test cases and GUI-based testing seem to be not adequately supported by existing AI-enhanced tools. Our findings are of interest to both professionals and researchers to develop innovative tools and technologies capable of supporting the identified solutions.
8.2 Observations from the Interviews